Need detailed explanation.

Linzer Augen are a delicious Austrian pastry many people like to consume around Christmas time. To Produce a Linzer Auge you need two plain biscuits and one spoon of jam. What is the minimum cost of producing 50000 Linzer Augen when one biscuit costs 2 cents and one spoon of jam costs 5 cents? Wählen Sie eine Antwort: O a. 45000 cents b. 35000 O c. 3500 cents d. 4500
Answer #1

d.4500

Number of biscuits required for 50000 Linzer Augen = 50000*biscuits required for 1 Linzer Augen=50000*2 = 100000

Number of spoons of jam required for 50000 Linzer Augen = 50000*spoons of jam required for 1 Linzer Augen=50000*1 = 50000

Cost of producing 50000 Linzer Augen = cost of 100000 biscuits+cost of 50000 spoons of jam = 2*100000+5*50000

=450000 cents = 450000/100 euro = 4500 euro